I am having trouble trying to graph the equation C = 6525 + 225S. The graph shown has C from 1000 to 9000 and S from 1 to 10. Where do the points go? Thanks so much if anuone can help me.
Answer by feliz1965(151)

I answer this question before.
I used the same equation listed here except for the notation.
I used C(S) = 6525 + 225(S).
Look, we let S equal the following values: 1, 2, 3, and 4. I selected these values but you can pick any positive numbers you like.
I will use your equation without the notation C(S), which is read "C of S."
C = 6525 + 225(1)
C = 6750
Our first point to graph is (1,6750).
C = 6525 + 225(2)
C = 6975
Our second point is (2,6975).
C = 6525 + 225(3)
C = 7200
Our this pointis (3,7200)
Lastly, we let S = 4 and do the same thing.
C = 6525 + 225(4)
C = 7425
Our fourth point is (4,7425)
All points have the form (x,y).
So, x represents the x-axis and y represents the y-axis.
Plot each point above on the coordinate plane.
If you have graph paper, that's even better.
Understand?